OpenStudy (graumann1128):

2(n - 3) + 5 = 3(n - 1) @quertty123

OpenStudy (kailey7277):

First, you would use the distributive property. To do that, multiply the 2 in front of the first set of parentheses by everything inside. \[2*n=2n\] \[2*-3=-6\] Then, you would combine those terms: \[2n-6\] Do the same thing with the parentheses on the other side of the equation: \[3*n=3n\] \[3*-1=-3\] Then, combine them: \[3n-3\] Don't forget about the -5 on the left side of the equation when you rewrite the whole thing after distributing. \[2n-6+5=3n-3\] You would combine the like terms, -6 and 5, on the left side. \[2n-1=3n-3\] Then, add 1 to both sides to get rid of the 1 on the left side. \[2n-1=3n-3\] \[+1..............+1\] \[2n=3n-2\] Then, subtract 3n from both sides. \[2n=3n-2\] \[-3n....-3n\] \[-n=-2\] Finally, divide both sides by negative one to make your variable positive. \[-n=-2\] \[\div-1..\div-1\] \[n=2\]